solve the equation 4x^2-6=10x, using FOIL.
Answer by jim_thompson5910(35256)

4x^2-6=10x


4x^2-6-10x=0


4x^2-10x-6=0


2(2x^2-5x-3)=0


2x^2-5x-3=0/2


2x^2-5x-3=0


(2x+1)(x-3)=0


2x+1=0 or x-3=0


x=-1/2 or x=3